the odds against the dispute manager x sitting the wage dispute with the workers are 8:6 and odds in favour of manager ysetting the same dispute is 14:16.
1. what is the chance that neither setting the dispute if both try independently.
2. what is the probability t6hat dispute is settled

Answer by stanbon(75887)

the odds against the dispute manager x settling the wage dispute with the workers are 8:6
P(against) = 8/14
P(for) = 6/14
------------------------
and odds in favor of manager y settling the same dispute is 14:16.
P(against) = 16/30
P(for) = 14/30
-------------------------
1. what is the chance that neither settling the dispute if both try independently.
Ans: (8/14)(16/30) = (4/7)(8/15) = 32/105
-------------------------------------------------
2. what is the probability that dispute is settled.
Ans: (6/14)(14/30) = (3/7)(7/15) = 1/5
==================
Cheers,
Stan H.
=================